# DiFSD: Ego-Centric Fully Sparse Paradigm with Uncertainty Denoising and Iterative Refinement for Efficient End-to-End Self-Driving

### v4

@@ Abstract
Current end-to-end autonomous driving methods resort to unifying modular designs for various tasks (e.g. perception, prediction and planning). Although optimized in a planning-oriented spirit with a fully differentiable framework, existing end-to-end driving systems without ego-centric designs still suffer from unsatisfactory performance and inferior efficiency, owing to the rasterized scene representation learning and redundant information transmission. In this paper, we revisit the human driving behavior and propose an ego-centric fully sparse paradigm, named DiFSD, for end-to-end self-driving. Specifically, DiFSD mainly consists of sparse perception, hierarchical interaction and iterative motion planner. The sparse perception module performs detection, tracking and online mapping based on sparse representation of the driving scene. The hierarchical interaction module aims to select the Closest In-Path Vehicle / Stationary (CIPV / CIPS) from coarse to fine, benefiting from an additional geometric prior. As for the iterative motion planner, both selected interactive agents and ego-vehicle are considered for joint motion prediction, where the output multi-modal ego-trajectories are optimized in an iterative fashion. Besides, both position-level motion diffusion and trajectory-level planning denoising are introduced for uncertainty modeling, thus facilitating the training stability and convergence of the whole framework. Extensive experiments conducted on nuScenes and Bench2Drive datasets demonstrate the superior planning performance and great efficiency of DiFSD.
@@ 1 Introduction
Literally, existing end-to-end models [11, 39, 17, 34] designed for reliable trajectory planning can be classified into two mainstreams as summarized in Fig. 1(a) and (b). The dense BEV-Centric paradigm [11, 39] performs perception, prediction and planning consecutively upon the shared BEV (Bird’s Eye View) features, which are computationally expensive leading to inferior efficiency. The sparse Query-Centric paradigm [34] utilizes sparse representation to achieve scene understanding and joint motion planning, thus improving the overall efficiency. However, object-intensive motion prediction inevitably causes computational redundancy and violates the driving habits of human drivers, who usually only concentrate on the Closest In-Path Vehicle / Stationary (CIPV / CIPS) which are more likely to affect the driving intention and trajectory planning of ego-vehicle. Meanwhile, excessive interaction with irrelevant agents will be conversely adverse to the ego-planning. Therefore, the planning performance remains unsatisfactory in both planning safety, comfort and personification.
@@
To this end, we propose DiFSD, an Ego-Centric fully sparse paradigm as shown in Fig. 1(c). Specifically, DiFSD mainly consists of sparse perception, hierarchical interaction and iterative motion planner. In the sparse perception module, multi-scale image features extracted from visual encoder are adopted for object detection, tracking and online mapping simultaneously in a sparse manner. Then the hierarchical interaction performs ego-centric and object-centric dual interaction to select the CIPV / CIPS with the help of an additional geometric prior. Thus the interactive queries can be selected gradually from coarse to fine. As for the motion planner, the mutual information between sparse interactive queries and ego-query is considered for motion prediction in a joint decoder, which is neglected in previous methods [11, 17] but is essential especially in the scenarios like intersections. To ensure the planning rationality and selection accuracy of interactive queries, the iterative planning optimization is further applied to the multi-modal proposal ego-trajectories, through continually updating the reference line and ego-query. Moreover, both position-level motion diffusion and trajectory-level planning denoising are introduced for stable training and fast convergence. It can not only model the uncertain positions of interactive agents for motion prediction, but also enhance the quality of trajectory refinement with arbitrary offsets. With above elaborate designs, DiFSD exhibits the great potential of fully sparse paradigm for end-to-end autonomous driving, which significantly reduces the average L2 error by 56% and collision rate by 92% than UniAD [11] respectively. Notably, our DiFSD-S achieves faster running efficiency as well. In sum, the main contributions of our work are as follows:
@@
We propose an ego-centric Fully Sparse paradigm for end-to-end self-Driving, named as DiFSD, without any computationally intensive dense scene representation learning and redundant environmental modeling, which is proven to be effective and efficient for ego-planning.
@@
We introduce a geometric prior through intention-guided attention, where the Closest In-Path Vehicle / Stationary (CIPV / CIPS) are gradually picked out through ego-centric cross attention and selection. Besides, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are adopted for uncertainty modeling of motion planning respectively.
@@ 2.1 End-to-End Perception
Recent years witness remarkable progress achieved in multi-view 3D detection, which mainly build elaborate designs upon the dense BEV (Bird’s Eye View) or sparse query features. To generate BEV features, LSS [32] lifts 2D image features to 3D space using depth estimation results, which are then splatted into BEV plane. Follow-up works apply such operation to perform view transform for 3D detection task [14, 13, 19, 7]. Differently, some works [20, 38, 12] project a series of predefined BEV queries in 3D space to the image domain for feature sampling. As for the sparse fashion, current methods [36, 26, 27, 25, 24] adopt a set of sparse queries to integrate spatial-temporal aggregations from multi-view image feature sequence for iterative anchor refinement.
2.2 Online Mapping
@@
Maps could provide important static scenario information to ensure driving safety. Current works [18, 28, 23, 40] manage to construct online maps with on-board sensors, instead of using HD-Map which is labor intensive and expensive. HDMapNet [18] achieves this aim through BEV semantic segmentation and heuristic post-processing to generate map instances. VectorMapNet [28] introduces a two-stage auto-regressive transformer to refine map elements consecutively. MapTR [23] regards map elements as a set of points with equivalent permutations, while StreamMapNet [40] adopts a temporal fusion strategy to enhance the performance. However, all of them reply on dense BEV features for online map construction, which is computationally intensive and not extensible to the sparse manner.
2.3 End-to-End Motion Prediction
@@
Motion prediction of surrounding agents in an end-to-end fashion can relieve the accumulative error between standalone models. FaF [31] predicts both current and future bounding boxes from images using a single convolution network. IntentNet [2] attempts to reason high-level behavior and long-term trajectories simultaneously. PnPNet [22] aggregate trajectory-level features for motion prediction through an online tracking module. ViP3D [6] takes images and HD-Map as input, and adopts agent queries to conduct tracking and prediction. PIP [16] further proposes to replace HD-Map with local vectorized map.
2.4 End-to-End Planning
@@
End-to-end planning paradigm either unites modules of perception and prediction [11, 17, 42, 39], or adopts a direct optimization on planning without intermediate tasks [3, 4, 33], which lack interpretability and are hard to optimize. Recently, UniAD [11] presents a planning-oriented model which integrates various tasks in the dense BEV-Centric paradigm, achieving convincing performance. VAD [17] learns vectorized scene representations and improves planning safety with explicit constraints. GraphAD [42] constructs the interaction scene graph to model both dynamic and static relations. SparseDrive [34] introduces the symmetric sparse perception for parallel motion planner. However, using straightforward designs and exhaustive modeling without ego-centric interaction, will inevitably lead to unsatisfactory planning performance and inferior efficiency.
@@ 3.1 Overview
The overall framework of proposed DiFSD is illustrated in Fig. 2, which deals with the end-to-end planning task in an ego-centric fully sparse paradigm. Specifically, DiFSD mainly consists of four parts: visual encoder, sparse perception, hierarchical interaction and iterative motion planner. First, the visual encoder extracts multi-scale spatial features from given multi-view images. Then the sparse perception takes the encoded features as input to perform detection, tracking and online mapping simultaneously. In the hierarchical interaction module, the ego query equipped with a geometric prior is introduced to pick out the interactive queries through ego-centric cross attention and hierarchical selection. In the iterative motion planner, both interactive agents and ego-vehicle are considered for joint motion prediction, then the predicted multi-modal ego-trajectories are further optimized iteratively. Meanwhile, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are conducted for uncertainty modeling of motion and planning tasks respectively.
@@ 3.3 Sparse Perception
After extracting the multi-view visual features from sensor images using the visual encoder [8], sparse query-based perception method proposed in [24, 25] is extended to perform unified detection and online mapping in parallel with the symmetric architecture as adopted in [34].
Detection. Following the previous methods [25, 24], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box can be denoted as:
which contains location, dimension, yaw angle as well as velocity respectively. Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through deformable aggregation of sample features projected from key points of . The updated instance features are adopted to predict the classification scores and box offsets respectively. Temporal instance denoising is introduced to improve model stability.
@@ 3.4 Ego-Env Hierarchical Interaction
Specifically, we use four MLPs to encode the ego-intention respectively, including velocity, acceleration, angular velocity and driving command. And then we concatenate these embeddings to obtain ego-intention features , which are further concatenated with the position embeddings of a group of pre-defined locations to cover densely distributed grid cells in the BEV plane. The position of each grid cell is represented as . Finally, the concatenated geometric features are fed to a single SE [9] block to learn response map , which is supervised by the normalized minimum distance from to the ego future waypoints. The motivation is that the Closest In-Path Vehicle / Stationary are prone to affect the ego-intention, and vice versa.
With the predicted response map , we first generate the reference line through row-wise thresholding, which are further used to generate the normalized distance map (See Fig. 2). Then we can obtain the geometric score for each surrounding query by referring to the . The reason why we don’t get the geometric score from directly is that the imbalanced distribution of ego-intention and future waypoints may lead to the inferior quality of .
@@
where the distance-prior is weighted with the attention score for both interaction and selection. is inner product, is dot product, and is the channel dimension.
Coarse-to-Fine Selection. To capture the interaction information from coarse to fine, we stack dual-interaction layers in a cascaded manner, where a top-K operation is appended between each two consecutive layers, thus the interactive objects can be gradually selected for latter prediction and planning usages. We claim that only a few interactive objects need to be considered for motion prediction, which are enough yet efficient for ego-centric path planning, instead of all detected agents existing in the driving scene.
@@ 3.5 Iterative Motion Planner
As shown in Fig. 2, the iterative motion planner is designed to conduct motion prediction for both interactive agents and ego-vehicle, and then optimize the proposal ego-trajectory with both safety and kinematic constrains iteratively.
Joint Motion Prediction. With regard to the trajectory prediction, both surrounding agents and ego-vehicle are adopted for motion modeling in a joint decoder, unlike previous works [11, 39, 17] which neglect the crucial interactions between near agents and ego-vehicle when making motion predictions, especially in the common scenarios like intersections. Another difference is that only the interactive objects (CIPV) sparsely selected in the former module are considered, instead of all detected agents in the driving scene which maybe irrelevant to the ego-vehicle planning. As for the joint motion decoder, we prepare three copies of ego query to indicate different driving intentions (i.e., turn left, turn right and keep forward), which are combined with to conduct agent-level self attention and agent-map cross attention respectively. And then we concatenate these output attended features to predict multi-modal trajectories , and classification scores , for both agents and ego-vehicle, where is the number of driving command for planning, are the mode number, are the future timestamps.
Planning Optimization. With the predicted multi-intention and multi-modal trajectories of ego-vehicle, we can select the most probable proposal trajectory with the input driving command and classification score . As shown in Fig. 3(b), ego-agent, ego-map and ego-navigator cross attentions are conducted consecutively for planning optimization, where the offsets for each future waypoint are predicted upon the proposal trajectory respectively with several planning constraints proposed in [17] to ensure safety.
Iterative Refinement. To further promote the stability and performance of the whole end-to-end system, an additional iterative refinement strategy is proposed to continuously update the reference line and distance map with refined ego trajectory as illustrated in Fig. 2, thus ensuring the interaction quality and selection accuracy of interactive queries.
@@ 3.6 Uncertainty Denoising
On one hand, position-level diffusion process is performed on ground-truth boxes of interactive agents for additional trajectory prediction of noisy agents:
@@
On the other hand, trajectory-level denoising process is also introduced for robust offset prediction of proposal trajectory of ego-vehicle in the planning optimization stage. Different from the position diffusion of detection or motion query described above, we apply the random noise to trajectory offsets of ego-vehicle , where depends on the Final Displacement (FD) of ground-truth ego future trajectory.
@@
| ST-P3 Metrics | ST-P3 [10] | EfficientNet-b4 | 1.33 | 2.11 | 2.90 | 2.11 | 0.23 | 0.62 | 1.27 | 0.71 | 628.3 | 1.6 |
| OccNet [35] | ResNet101-DCN | 1.29 | 2.13 | 2.99 | 2.13 | 0.21 | 0.59 | 1.37 | 0.72 | - | - | |
| FusionAD∗ [39] | ResNet101+SECOND [37] | - | - | - | 1.03 | 0.25 | 0.13 | 0.25 | 0.21 | - | - | |
@@
| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.03 | 0.07 | 0.21 | 0.10 | 67.7 | 14.8 | |
| DiFSD-S (PV) | ResNet50 | 0.15 | 0.31 | 0.56 | 0.33 | 0.00 | 0.06 | 0.19 | 0.08 | 93.7 | 10.7 | |
| DiFSD-B (PV) | ResNet101 | 0.15 | 0.30 | 0.54 | 0.32 | 0.00 | 0.04 | 0.15 | 0.06 | 119.6 | 8.4 | |

### v5

# EgoFSD: Ego-Centric Fully Sparse Paradigm with Uncertainty Denoising and Iterative Refinement for Efficient End-to-End Self-Driving

@@ Abstract
Current End-to-End Autonomous Driving (E2E-AD) methods resort to unifying modular designs for various tasks (e.g. perception, prediction and planning). Although optimized with a fully differentiable framework in a planning-oriented manner, existing end-to-end driving systems lacking ego-centric designs still suffer from unsatisfactory performance and inferior efficiency, due to rasterized scene representation learning and redundant information transmission. In this paper, we propose an ego-centric fully sparse paradigm, named EgoFSD, for end-to-end self-driving. Specifically, EgoFSD consists of sparse perception, hierarchical interaction and iterative motion planner. The sparse perception module performs detection and online mapping based on sparse representation of the driving scene. The hierarchical interaction module aims to select the Closest In-Path Vehicle / Stationary (CIPV / CIPS) from coarse to fine, benefiting from an additional geometric prior. As for the iterative motion planner, both selected interactive agents and ego-vehicle are considered for joint motion prediction, where the output multi-modal ego-trajectories are optimized in an iterative fashion. In addition, position-level motion diffusion and trajectory-level planning denoising are introduced for uncertainty modeling, thereby enhancing the training stability and convergence speed. Extensive experiments are conducted on nuScenes and Bench2Drive datasets, which significantly reduces the average L2 error by 59% and collision rate by 92% than UniAD while achieves 6.9 faster running efficiency.
@@ 1 Introduction
Literally, existing end-to-end models [12, 47, 17, 39, 44, 45, 46] designed for reliable trajectory planning can be classified into two mainstreams as summarized in Fig. 1(a) and (b). The dense BEV-Centric paradigm [12, 47] performs perception, prediction and planning consecutively upon the shared BEV (Bird’s Eye View) features, which are computationally expensive leading to inferior efficiency. The sparse Query-Centric paradigm [39] utilizes sparse representation to achieve scene understanding and joint motion planning, thus improving the overall efficiency. However, object-intensive motion prediction inevitably causes computational redundancy and violates the driving habits of human drivers, who usually only concentrate on the Closest In-Path Vehicle / Stationary (CIPV / CIPS) which are more likely to affect the driving intention and trajectory planning of ego-vehicle. Meanwhile, excessive interaction with irrelevant agents will be conversely adverse to the ego-planning. Therefore, the planning performance remains unsatisfactory in both planning safety, comfort and personification.
@@
To this end, we propose EgoFSD, an Ego-Centric fully sparse paradigm as shown in Fig. 1(c). Specifically, EgoFSD consists of sparse perception, hierarchical interaction and iterative motion planner. In the sparse perception module, multi-scale image features extracted from visual encoder are adopted for object detection and online mapping simultaneously in a sparse manner. Then the hierarchical interaction performs ego-centric and object-centric dual interaction to select the CIPV / CIPS with the help of an additional geometric prior. Thus the interactive queries can be selected gradually from coarse to fine. As for the motion planner, the mutual information between sparse interactive queries and ego-query is considered for motion prediction in a joint decoder, which is neglected in previous methods [12, 17] but is essential especially in the scenarios like intersections. To ensure the planning rationality and selection accuracy of interactive queries, the iterative planning optimization is further applied to the multi-modal proposal ego-trajectories, through continually updating the reference line and ego-query. Moreover, both position-level motion diffusion and trajectory-level planning denoising are introduced for stable training and fast convergence. It can not only model the uncertain positions of interactive agents for motion prediction, but also enhance the quality of trajectory refinement with arbitrary offsets. With above elaborate designs, EgoFSD exhibits the great potential of fully sparse paradigm for end-to-end autonomous driving, which significantly reduces the average L2 error by 59% and collision rate by 92% than UniAD [12] respectively. Notably, our EgoFSD-B achieves faster running efficiency as well. In sum, the main contributions of our work are as follows:
@@
We propose an Ego-Centric Fully Sparse paradigm for end-to-end self-Driving, named EgoFSD, without dense representation learning and redundant environmental modeling, which is proven to be effective for ego-planning.
@@
We introduce a geometric prior through intention-guided attention, where the Closest In-Path Vehicle / Stationary (CIPV / CIPS) are gradually picked out through ego-centric cross attention and selection. Besides, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are adopted for uncertainty motion modeling.
@@ 2 Related Work
Object Detection. Recent years witness remarkable progress achieved in multi-view 3D detection, which mainly build elaborate designs upon the dense BEV (Bird’s Eye View) or sparse query features. To generate BEV features, LSS [36] lifts 2D image features to 3D space using depth estimation results, which are then splatted into BEV plane. Follow-up works apply such operation to perform view transform for 3D detection task [14, 21, 7]. Differently, some works [22, 43, 13] project a series of predefined BEV queries in 3D space to the image domain for feature sampling. As for the sparse fashion, current methods [42, 29, 30, 28, 27] adopt a set of sparse queries to integrate spatial-temporal aggregations from multi-view image feature sequence for iterative anchor refinement.
@@
Online Mapping. Maps could provide important static scenario information to ensure driving safety. Current works [20, 31, 25, 48] manage to construct online maps with on-board sensors, instead of using HD-Map which is labor intensive and expensive. HDMapNet [20] achieves this aim through BEV semantic segmentation and heuristic post-processing to generate map instances. VectorMapNet [31] introduces a two-stage auto-regressive transformer to refine map elements consecutively. MapTR [25] regards map elements as a set of points with equivalent permutations, while StreamMapNet [48] adopts a temporal fusion strategy to enhance the performance. However, all of them rely on dense BEV features for online map construction, which is computationally intensive and thus inefficient.
@@
Motion Prediction. Predicting agent future trajectories is essential for the autonomous vehicle to understand motion intention of surrounding agents. FaF [34] predicts both current and future boxes from images using a single deep network. IntentNet [2] attempts to reason high-level behavior and long-term trajectories simultaneously. PnPNet [24] aggregate trajectory-level features for motion prediction through an online tracking module. ViP3D [6] takes images and HD-Map as input, and adopts agent queries to conduct tracking and prediction. PIP [16] further proposes to replace HD-Map with local vectorized map.
@@
End-to-End Planning. End-to-end planning paradigm either unites modules of perception and prediction [17, 51, 47, 35, 37], or adopts a direct optimization on planning without intermediate tasks [3, 4, 19], which lack interpretability and are hard to optimize. Recently, UniAD [12] presents a planning-oriented model which integrates various tasks in the dense BEV-Centric paradigm, achieving convincing performance. VAD [17] learns vectorized scene representations and improves planning safety with explicit constraints. GraphAD [51] constructs the interaction scene graph to model both dynamic and static relations. SparseDrive [39] introduces the symmetric sparse perception for parallel motion planner, which consumes more computation cost due to the repeated query projection and deformable feature aggregation, without a shared 3D feature [12, 17]. Besides, using straightforward designs and exhaustive modeling without ego-centric interaction, will inevitably lead to unsatisfactory planning performance and inferior efficiency.
@@ 3.1 Overview
The overall framework of proposed EgoFSD is illustrated in Fig. 2, which deals with the end-to-end planning task in an ego-centric fully sparse paradigm. Specifically, EgoFSD mainly consists of four parts: visual encoder, sparse perception, hierarchical interaction and iterative motion planner. First, the visual encoder extracts multi-scale spatial features from given multi-view images. Then the sparse perception takes the encoded features as input to perform detection, tracking and online mapping simultaneously. In the hierarchical interaction module, the ego query equipped with a geometric prior is introduced to pick out the interactive queries through ego-centric cross attention and hierarchical selection. In the iterative motion planner, both interactive agents and ego-vehicle are considered for joint motion prediction, then the predicted multi-modal ego-trajectories are further optimized iteratively. Meanwhile, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are conducted for uncertainty modeling of motion and planning tasks respectively.
@@ 3.3 Sparse Perception
After extracting the multi-view visual features from sensor images using the visual encoder [8], sparse query-based perception method proposed in [27, 29, 40] is extended to perform unified detection and online mapping in parallel with the symmetric architecture as adopted in [39].
Detection. Following [27, 28], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box can be denoted:
which contains location, dimension, yaw angle as well as velocity respectively. Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through iterative decoding. The updated instance features are adopted to predict the classification scores and box offsets respectively. Refer to [40], hybrid attention between current propagated queries and history memory queue is conducted for temporal modeling. Besides, temporal instance denoising is introduced to improve model stability.
@@ 3.4 Ego-Env Hierarchical Interaction
Specifically, we use four MLPs to encode the ego-intention respectively, including velocity, acceleration, angular velocity and driving command. Then we concatenate these embeddings to obtain ego-intention features , which are further concatenated with the position embeddings of a group of pre-defined locations to cover densely distributed grid cells in the BEV plane. Finally, the concatenated geometric features are fed to a single SE [9] block to learn response map , which is supervised by the normalized minimum distance from to interpolated future trajectory with waypoints. Thus, for each grid , the response value can be formulated as:
where is set to 30 for trajectory interpolation empirically. The motivation is that the Closest In-Path Vehicle / Stationary are prone to affect the ego-intention, and vice versa.
With the predicted , we first generate the reference line through row-wise thresholding, which are further used to generate the normalized distance map (See Fig. 2). Then we can obtain the geometric score for each object query by referring to the . The reason why we don’t get the geometric score from directly is that the imbalanced distribution of ego-intention and future waypoints [23] may lead to the inferior quality of .
@@
where the distance-prior is weighted with the attention score for both interaction and selection. is inner product, is Hadamard product, and is the channel dimension.
Coarse-to-Fine Selection. To capture the interaction information from coarse to fine, we stack dual-interaction layers in a cascaded manner, where a top- operation is appended between each two consecutive layers, thus the interactive objects can be gradually selected for latter prediction and planning usages. We claim that only a few interactive objects need to be considered for motion prediction, which are enough yet efficient for ego-centric path planning, instead of all detected agents existing in the driving scene. is set to 3 and empirically.
@@ 3.5 Iterative Motion Planner
As shown in Fig. 2, the iterative motion planner is designed to conduct motion prediction for both interactive agents and ego-vehicle, and then optimize the ego-trajectory with both safety and kinematic constrains iteratively.
Joint Motion Prediction. With regard to the trajectory prediction, both surrounding agents and ego-vehicle are adopted for motion modeling in a joint decoder, unlike previous works [12, 17, 47] which neglect the crucial interactions between near agents and ego-vehicle when making motion predictions, especially in the common scenarios like intersections. Another difference is that only the interactive objects (CIPV) sparsely selected in the former module are considered, instead of all detected agents in the driving scene which maybe irrelevant to the ego-vehicle planning. As for the joint motion decoder, we prepare three copies of ego query to indicate different driving intentions (i.e., turn left, turn right and keep forward), which are combined with to conduct agent-level self-attention and agent-map cross attention respectively. And then we concatenate these enhanced features to predict multi-modal trajectories , and classification scores , for both agents and ego-vehicle, where is the number of driving command for planning, are the mode number, are the future timestamps.
Planning Optimization. With the predicted multi-intention and multi-modal trajectories of ego-vehicle, we can select the most probable proposal trajectory with the input driving command and classification score . As shown in Fig. 3(b), ego-agent, ego-map and ego-navigator cross attentions are conducted consecutively for planning optimization. And the offsets for each future waypoint are predicted upon the proposal trajectory respectively with several planning constraints proposed in [17] to ensure safety.
Iterative Refinement. To ensure the interaction quality and selection accuracy of interactive queries, an additional iterative refinement strategy is proposed to continuously update the reference line and distance map with refined ego trajectory as illustrated in Fig. 2.
@@ 3.6 Uncertainty Denoising
On one hand, position-level diffusion process is performed on Top- boxes of interactive agents for additional trajectory prediction of noisy agents:
@@
On the other hand, trajectory-level denoising process is also introduced for robust offset prediction of proposal trajectory of ego-vehicle in the planning optimization stage. Different from the position diffusion of agent query on the purpose of detection and motion, we apply the random noise to the ground-truth trajectory of ego-vehicle:
where indicate noisy trajectory proposals and the noise scale depends on the Final Displacement (FD) of .
@@
| ST-P3 Metrics | FF∗ [10] | - | 0.55 | 1.20 | 2.54 | 1.43 | 0.06 | 0.17 | 1.07 | 0.43 | - | - |
| EO∗ [18] | - | 0.67 | 1.36 | 2.78 | 1.60 | 0.04 | 0.09 | 0.88 | 0.33 | - | - | |
| ST-P3 [11] | EfficientNet-b4 | 1.33 | 2.11 | 2.90 | 2.11 | 0.23 | 0.62 | 1.27 | 0.71 | 628.3 | 1.6 | |
@@
| EgoFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.03 | 0.07 | 0.21 | 0.10 | 67.7 | 14.8 | |
| EgoFSD-S (PV) | ResNet50 | 0.15 | 0.31 | 0.56 | 0.33 | 0.00 | 0.06 | 0.19 | 0.08 | 59.8 | 16.7 | |
| EgoFSD-B (PV) | ResNet101 | 0.12 | 0.28 | 0.52 | 0.30 | 0.00 | 0.04 | 0.15 | 0.06 | 80.5 | 12.4 | |